# Surface to Seafloor: A Generative Machine Learning Framework for Decoding the Ocean Interior State

### v1

@@ 1 Introduction
Importantly, probabilistic modeling allows for uncertainty quantification (UQ) (Forget et al., 2015; Souza et al., 2020): not only can samples from the inferred distribution be explored, but these samples also provide estimates of plausible subsurface fields when surface information is insufficient to constrain the output fully. Generative modeling offers a flexible approach to sampling from such distributions, with score-based diffusion models emerging as a particularly effective tool for high-dimensional statistical inference. Generative-adversarial models (GANs) have also been applied to predict ocean temperature profiles from sea surface temperature (Meng et al., 2021; Zhang et al., 2023), but these works did not leverage the potential for uncertainty quantification. More recently, Martin et al. (2025) have leveraged score-based generative methods in the context of ocean surface data assimilation but again, they did not focus on uncertainty quantification. Beyond conditional sampling, score-based generative models have demonstrated utility in diverse applications, including calculating response functions (Giorgini et al., 2024), estimating the dimensionality of data manifolds (Stanczuk et al., 2023), stabilizing the dynamical evolution neural network models (Pedersen et al., 2025), or devising stochastic analogs to chaotic equations (Giorgini et al., 2025).
@@ 2.1 Ocean Simulation
Using Oceananigans.jl (Ramadhan et al., 2020; Silvestri et al., 2024a, b; Wagner et al., 2025), we solve the Boussinesq equations under the hydrostatic approximation. The dynamical core uses a nonlinear free surface and $z^{\star}$ coordinates. The prognostic variables are the horizontal velocities, $U$ and $V$, the SSH $\eta$, and the temperature $T$. A linear equation of state relates the buoyancy of seawater $B$ to the temperature, (Roquet et al., 2015). The figures will refer to the Oceananigans simulation output as “OcS” for brevity.
The idealized ocean domain extends from 15 to 75 degrees north in latitude and from 0 to 60 degrees in longitude with walls on all sides. The domain is 1800 m deep with a flat bottom. A zonal wind stress is applied at the surface, independent of longitude, with a cosine profile peaking eastward (positive) in the middle of the domain and at a minimum and westward (i.e. negative) at the northern and southern walls. The surface temperature is restored to a linear profile in latitude ranging from $30$ C on the southern boundary at $15^{\circ}$, to 00 C at the northern boundary at $75^{\circ}$. A quadratic bottom drag is imposed at the seafloor.
@@ 2.2 Score-Based Diffusion Model
where $\mathcal{P}_{M}$ is the marginal distribution of the subsurface fields. Most inference tasks lie between these two extremes, Equations 3 and 4.
@@ 3 Results
Given the ocean simulation data (OcS) from Section 22.1 and the neural network architecture from Section 22.2, we perform statistical inference by drawing 100 sample fields using the reverse diffusion process, given SSH. We test the enural network by feeding SSH from the last 50 years of the double-gyre simulation which were not used during training. The predictions of the subsurface variables are tested for the various caorse-grained SSH input data.
@@ 3.3 Heat Flux Predictions and Uncertainties with Degraded Surface Height Input
Figure 7: Zonal Average of Vertical and Horizontal Eddy Heat Flux for a Fixed Sea Surface Height at Various Depths. We show the zonal average of the vertical (top row) and meridional eddy heat flux (bottom row) over the domain for each latitude at a fixed moment time and various depths in blue. The neural network results using the full-resolution sea surface height for input are in red, with the solid red line representing the ensemble average heat flux. The ribbons represent the $0.6,0.7,0.8$, and $0.9$ quantiles, from darker to lighter, as computed with the 100-member ensemble. Each column shows a different depth, and the relative uncertainty (red ribbon region) increases as we go deeper into the domain.
Figure 8: Zonal Average of Vertical and Horizontal Eddy Heat Flux for a Fixed Sea Surface Height at a 400 Meter Depth. We show the zonal average of the vertical (top row) and meridional heat flux (bottom row) over the first fifteen degrees of the domain for each latitude at a fixed moment time and a depth of 400 meters in blue. The neural network results are in red, with the solid red line representing the ensemble average heat flux. The ribbons represent the $0.6,0.7,0.8$, and $0.9$ quantiles, from darker to lighter. Each column uses different levels of coarse-grained sea surface height. As the amount of coarse-graining increases, the confidence bounds (as represented by the red region) increase and reduce to the marginal distribution of the data distribution). Information about the location of the meander is lost when going from a factor of 8 reduction to a factor of 16 in reduction of resolution. See the first column and third row of Figure 4 for the corresponding SSH.
@@ 4 Conclusions
Our results demonstrate that generative models might capture the ocean interior’s statistical variability, even when conditioned on coarse-resolution surface data. However, decreasing surface observations’ spatial resolution or increasing inference depth diminishes predictive skill. In the limiting case, when the surface signal is insufficient to constrain the interior, the generative model reverts to sampling from the marginal distribution of the training data. As shown in Figure 6, the most pronounced deterioration in performance occurs when the surface resolution is degraded from 100km (coarsening factor $2^{2}$) to 200km ( $2^{3}$), particularly for the horizontal velocity components where the eddy length scale is on the order of 100km. For temperature, predictive accuracy declines when large-scale gradient information is lost. These findings suggest that improvements in satellite resolution—from current AVISO-class missions (200km) to SWOT (50km)—could significantly enhance our ability to infer subsurface ocean dynamics Ballarotta et al. (2019); Fu et al. (2024). Incorporating additional dynamical priors, such as quasi-geostrophic constraints or hybrid physics-informed machine learning architectures, may further improve the fidelity and robustness of generative reconstructions.

### v2

# Surface to Seafloor: A Generative AI Framework for Decoding the Ocean Interior State

@@ 1 Introduction
Importantly, probabilistic modeling allows for uncertainty quantification (UQ) (Forget et al., 2015; Souza et al., 2020): not only can samples from the inferred distribution be explored, but these samples also provide estimates of plausible subsurface fields when surface information is insufficient to constrain the output fully. Generative modeling offers a flexible approach to sampling from such distributions, with score-based diffusion models emerging as a particularly effective tool for high-dimensional statistical inference. Generative-adversarial models (GANs) have also been applied to predict ocean temperature profiles from sea surface temperature (Meng et al., 2021; Zhang et al., 2023), but these works did not leverage the potential for uncertainty quantification. More recently, Martin et al. (2025) have leveraged score-based generative methods in the context of ocean surface data assimilation but again, they did not focus on uncertainty quantification. Beyond conditional sampling, score-based generative models have demonstrated utility in diverse applications, including calculating response functions (Giorgini et al., 2024, 2025b), estimating the dimensionality of data manifolds (Stanczuk et al., 2023), stabilizing the dynamical evolution neural network models (Pedersen et al., 2025), or devising stochastic analogs to chaotic equations (Giorgini et al., 2025a).
@@ 2.1 Ocean Simulation
Using Oceananigans.jl (Ramadhan et al., 2020; Silvestri et al., 2024a; Wagner et al., 2025a; Silvestri et al., 2024b; Wagner et al., 2025b), we solve the Boussinesq equations under the hydrostatic approximation. The dynamical core uses a nonlinear free surface and $z^{\star}$ coordinates. The prognostic variables are the horizontal velocities, $U$ and $V$, the SSH $\eta$, and the temperature $T$. A linear equation of state relates the buoyancy of seawater $B$ to the temperature, (Roquet et al., 2015). The figures will refer to the Oceananigans simulation output as “OcS” for brevity.
The idealized ocean domain extends from 15 to 75 degrees north in latitude and from 0 to 60 degrees in longitude with walls on all sides. The domain is 1800 m deep with a flat bottom. A zonal wind stress is applied at the surface, independent of longitude, with a cosine profile peaking eastward (positive) in the middle of the domain at a value of 0.1 $0.1\text{ N}\text{m}^{-2}$ and at a minimum and westward (i.e. negative) at the northern and southern walls. The surface temperature is restored to a linear profile in latitude ranging from $30$ C on the southern boundary at $15^{\circ}$, to 00 C at the northern boundary at $75^{\circ}$. A quadratic bottom drag is imposed at the seafloor and convective adjustment is used for vertical mixing. The (dimensionless) quadratic drag coefficient is $10^{-3}$ and we use a horizontal diffusivity and viscosity of $2.5\times 10^{2}\text{ m}^{2}\text{s}^{-1}$, a background vertical diffusivity of $10^{-5}\text{ m}^{2}\text{s}^{-1}$, and background vertical viscosity of $10^{-2}\text{ m}^{2}\text{s}^{-1}$.
@@ 2.2 Score-Based Diffusion Model
where $\mathcal{P}_{M}$ is the marginal distribution of the subsurface fields. Most inference tasks lie between these two extremes, Equations 3 and 4. We comment that we use the word “marginal” as is commonly used in the statistics literature with regards to distributions.
@@ 3 Results
Given the ocean simulation data (OcS) from Section 22.1 and the neural network architecture from Section 22.2, we perform statistical inference by drawing 100 sample fields using the reverse diffusion process, given SSH. We test the enural network by feeding SSH from the last 50 years of the double-gyre simulation which were not used during training. The predictions of the subsurface variables are tested for the various coarse-grained SSH input data.
@@ 3.3 Heat Flux Predictions and Uncertainties with Degraded Surface Height Input
Figure 7: Zonal average of vertical and horizontal eddy heat flux for a fixed SSH at various depths. We show the zonal average of the vertical (top row) and meridional eddy heat flux (bottom row) over the domain for each latitude at a fixed moment time and various depths in blue. The neural network results using the full-resolution sea surface height for input are in red, with the solid red line representing the ensemble average heat flux. The ribbons represent the $0.6,0.7,0.8$, and $0.9$ quantiles, from darker to lighter, as computed with the 100-member ensemble. Each column shows a different depth, and the relative uncertainty (red ribbon region) increases as we go deeper into the domain.
Figure 8: Zonal average of vertical and horizontal eddy heat flux for a fixed SSH at a 400 meter depth. We show the zonal average of the vertical (top row) and meridional heat flux (bottom row) over the first fifteen degrees of the domain for each latitude at a fixed moment time and a depth of 400 meters in blue. The neural network results are in red, with the solid red line representing the ensemble average heat flux. The ribbons represent the $0.6,0.7,0.8$, and $0.9$ quantiles, from darker to lighter. Each column uses different levels of coarse-grained sea surface height. As the amount of coarse-graining increases, the confidence bounds (as represented by the red region) increase and reduce to the marginal distribution of the data distribution). Information about the location of the meander is lost when going from a factor of 8 reduction to a factor of 16 in reduction of resolution. See the first column and third row of Figure 4 for the corresponding SSH.
@@ 4 Conclusions
Our results demonstrate that generative models might capture the ocean interior’s statistical variability, even when conditioned on coarse-resolution surface data. However, decreasing surface observations’ spatial resolution or increasing inference depth diminishes predictive skill. In the limiting case, when the surface signal is insufficient to constrain the interior, the generative model reverts to sampling from the marginal distribution of the training data. As shown in Figure 6, the most pronounced deterioration in performance occurs when the surface resolution is degraded from 100km (coarsening factor $2^{2}$) to 200km ( $2^{3}$), particularly for the horizontal velocity components where the eddy length scale is on the order of 100km. For temperature, predictive accuracy declines when large-scale gradient information is lost. These findings suggest that improvements in satellite resolution—from current AVISO-class missions (200km) to SWOT (50km)—could significantly enhance our ability to infer subsurface ocean dynamics (Ballarotta et al., 2019; Fu et al., 2024). Incorporating additional dynamical priors, such as quasi-geostrophic constraints or hybrid physics-informed machine learning architectures, may further improve the fidelity and robustness of generative reconstructions.